# Demonstration of logical qubits and repeated error correction with better-than-physical error rates

### v2

@@ IV.1 Carbon code circuits
In the spirit of Ref. 22, instead of simply showing individual logical circuits that outperform their physical counterparts, we would like to show that families of circuits composed to perform some computation benefit from encoding. A crucial element of such a demonstration is repeated fault-tolerant error correction.
@@
The Carbon code, on the other hand, has a high threshold under the usual assumptions about parallelism, access to fresh ancillas, uniformity of the noise [49]. This proposal relies on teleportation-based syndrome extraction to achieve this performance [50]. Such an approach requires access to 3 code blocks within the hardware (one block for the data, and two blocks for the maximally entangled resource state).
Since the Carbon code requires 12 qubits per block, three blocks of the Carbon code would not fit in the 32-qubit H2 quantum processor [21]. Instead, we use a more compact scheme similar to Steane syndrome extraction [51], but derived from Knill’s approach to syndrome extraction. The key observation made in Ref. 49 is illustrated in Fig. 5, resulting in a circuit that essentially serializes the syndrome extracting teleportation into two 1-bit teleportations [52]—the only requirement being that one must be able to prepare the encoded |$\ket{0}$ and $\ket{+}$ states, apply an encoded $\rm CNOT$, and measure in the encoded $X$ and $Z$ eigenbases (and all of these operations are available in the Carbon code).
[…]
Figure 5: Syndrome information can be obtained by performing teleportation at the logical level, as described by Knill [50]. Taking the realization of the original teleportation circuit for two logical qubits encoded in a Carbon block, which required 3 encoded blocks (left), it is possible to rearrange commuting circuit components to arrive at a circuit that uses a sequence of two 1-bit teleportations [52] to extract syndrome information requiring only 2 encoded blocks at any given time (right) [49].
@@ IV.2 Physical baselines
Aside from syndrome extraction, we would like to demonstrate encoded gates acting on the encoded information. Using the resources available in the experiment, we may only consider operations acting on a single data block, as the ancilla block and additional physical ancillas used in state preparation (31 qubits in total) take up most of the remaining qubit resources available in H2.
Within a single block, the simplest fault-tolerant operations available for the Carbon code are the ${\rm H}\otimes{\rm H}$ gates within a block, and a $\rm CNOT$ between the two logical qubits within the block—the $\rm H$ is implemented via transversal $\rm H$ application followed by a qubit permutation, while the $\rm CNOT$ is implemented via a qubit permutation [53, 54]. Since qubit permutations can be applied in the H2 system without interaction between qubits, the permutations do not require additional ancillas to be fault-tolerant.
Combining gate teleportation and permutation-based gates allows to perform logical circuits of the form depicted in Fig. 6 in a single round of error correction. The $\rm CNOT$s and $\rm SWAP$s of logical qubits within a block and their composition can be implemented by qubit permutations.
[…]
Figure 6: Logical circuit family that can be implemented in a modified single round of error correction with the $[[12,2,4]]$ Carbon code, assuming $(\alpha,\beta,\delta,\gamma)\in\{0,\pm\frac{\pi}{2},\pm\pi\}$. The logical single-qubit rotations require modified state preparation during syndrome extraction, and the order of $X$ and $Z$ rotations can be transposed by modifying the order of the 1-bit teleportations, but neither of these modification impact the ability to extract syndrome information. Any combination of gates in the dashed boxes can be implemented implemented by a permutation of the physical qubits.
@@
The comparison between a sequence of two $\rm CNOT$s and a round of error correction is motivated by the proposal in Ref. 22. Instead of benchmarking the entire circuit family implied by Fig. 6, we choose to compare against circuits two $\rm CNOT$s since these are the physical gates with highest error rates in the physical system. Although in a fully scalable sub-threshold setting all logical gates can be made to improve upon their physical counterparts (by going to sufficiently high distance), the first bar that an experimental demonstration must meet is the demonstration of some non-trivial logical circuit that improves over its physical counterpart, which is our aim here.
@@ IV.3 Compiler optimizations
Time spent shuttling and cooling ions can lead to significant memory error, especially in error correction circuits that tend to leave some qubits idling for extended times [12]. Effective memory errors in H2 were previously reported in Ref. 21 for certain types of circuits with random connectivity, but the circuits used in this study are highly structured, leading to more coherent accumulation of noise, which is especially difficult to model accurately. To mitigate this problem, for the Carbon code experiments, we (1) modified the optimization of transport operations in the compiler and (2) incorporated dynamical decoupling pulses into the compiled circuits.
@@ IV.4 Experimental results
We observe a gain of more than an order of magnitude for a single round of error correction, as evidenced by a logical error rate of $0.03\%$, while the two physical baselines yield error rates of $0.51\%$ (for physical teleportations) and $0.35\%$ (a sequence of 2 $\rm CNOT$s). For two rounds of error correction, the logical error rates increase to $0.4\%$, while the two physical baselines have error rates increase to $0.9\%$ and $0.8\%$ respectively, which represents a logical gain of 2 over the physical baselines. For three rounds of error correction the uncertainty in the logical error rate estimates do now allow us to conclusively point to an improvement (the error bars for the logical error rate estimate the and physical baseline error rates overlap slightly due to a limitation in the number of runs). Details for each experiment and associated statistical uncertainties can be found in Table 3, and the results are illustrated in Fig. 7.
Although the gains over the physical error rates appear to behave non-linearly, the trend for the observed error rate per round of error correction is consistent with the expected linear behavior for a small number of repetitions, as illustrated in Fig. 7. The maximum-likelihood fits yield error rates per round of the teleportation and the $\rm CNOT$ baseline of $0.42\%\pm 0.09\%$ and $0.41\%\pm 0.08\%$ respectively, while the fit for each round of error correction points to $0.41\%\pm 0.13\%$—indicating that the differences in the error rates per round are not statistically significant.
The apparent non-linear trend in the gain is an artifact of the different $y$-intercepts for the linear trends of each of the sets of experiments. For the physical experiments, the $y$-intercept is set by the physical state-preparation and measurement (SPAM) error rate. This is also the case for the logical circuits, but details of the circuit implementation leads to a negative $y$-intercept despite positive SPAM error rates. This can be explained by considering optimizations in scheduling of the operations in the experimental device. In particular, since memory and transport errors are the dominating source of error in the experiments (see Section II), we estimate the optimized compiled circuit to have roughly $12r-$ 1 periods of significant waiting and transport for $r$ rounds of error correction (due details of the initial state preparation and final state measurement), so the logical error rate for $r$ rounds is estimated to be roughly
@@
The pre- and post-rejection rates also serve to inform whether the error correction circuits are behaving as expected. In particular we expect the pre- and post-rejection rates to be linear in the number of repetitions (to leading order), and the experimental data is highly consistent with that prediction (see Appendix B).
@@
1 round unencoded baseline (2 teleports) 20,000 102 $0.51\%_{-0.09\%}^{+0.11\%}$
unencoded baseline (2 $\rm CNOT$s) 20,000 70 $0.35\%_{-0.08\%}^{+0.09\%}$
encoded, pre- and post-selection 10,000 7,054 7,022 1432 2 $0.03\%_{-0.03\%}^{+0.06\%}$ 12–17
2 rounds unencoded baseline (4 teleports) 20,000 189 $0.9\%_{-0.1\%}^{+0.1\%}$
unencoded baseline (4 $\rm CNOT$s) 20,000 154 $0.8\%_{-0.1\%}^{+0.1\%}$
encoded, pre- and post-selection 14,548 7,685 7,345 2,948 28 $0.4\%_{-0.1\%}^{+0.2\%}$ 2–2.25
3 rounds unencoded baseline (6 teleports) 20,000 271 $1.4\%_{-0.2\%}^{+0.2\%}$
unencoded baseline (6 $\rm CNOT$s) 20,000 236 $1.2\%_{-0.1\%}^{+0.2\%}$
encoded, pre- and post-selection 10,000 4,010 3,645 1,953 28 $0.8\%_{-0.3\%}^{+0.3\%}$ 1.5–1.75
@@
Figure 7: Observed error rate for circuits with 1 to 3 rounds of error correction with the $[[12,2,4]]$ Carbon code (green circles) and physical baselines (blue diamond for pairs of 1-bit teleportations, and orange squares for pairs of $\rm CNOT$s). Results are offset along the x-axis for clarity. Linear fits are obtained by maximum-likelihood estimation (see Appendix A for details).
@@ V Summary and Outlook
We have demonstrated several fault-tolerant circuits outperforming their physical counterparts in a state-of-the-art QCCD trapped-ion system. Using different codes and protocols, we demonstrated error rates 4.7 to 800 times lower than the physical error rates in Bell correlation experiments with the $[[7,1,3]]$ and the $[[12,2,4]]$ codes. Moreover, we demonstrated up to 3 rounds of error correction for the $[[12,2,4]]$ quantum code by combining error correction and error rejection, with logical error rates below a physical baseline circuit of two $\rm CNOT$s per round. The error rate per round of error correction is consistent with error rates that are comparable with the physical error rates of two $\rm CNOT$s, and potentially better, although more data would be necessary to state this conclusively. With these results, we have demonstrated that current quantum processors are already able to reduce error rates in small circuits through quantum error correction. Future work will focus on extending these results to a richer set of fault-tolerant logical gadgets, and to ultimately enable universal fault-tolerant quantum computation while continuing to lower the achievable logical error rates. A significant milestone will be to demonstrate a universal family of quantum circuits with logical error rates approaching $10^{-8}$.
@@ Acknowledgements.
We thank the broader quantum teams at Microsoft Azure Quantum and Quantinuum for many helpful discussions. We thank Jeongwan Haah for insightful discussions and helpful feedback, and Adam Paetznick for his contributions to the discussions and assitance in designing the Carbon code experiments. A special thanks to Z. Alam for making this collaboration possible.
Appendix A Statistical analysis and error bars

### v3

@@ IV.1 Carbon code circuits
In the spirit of Ref. 23, instead of simply showing individual logical circuits that outperform their physical counterparts, we would like to show that families of circuits composed to perform some computation benefit from encoding. A crucial element of such a demonstration is repeated fault-tolerant error correction.
@@
The Carbon code, on the other hand, has a high threshold under the usual assumptions about parallelism, access to fresh ancillas, uniformity of the noise [42, 43]. This proposal relies on teleportation-based syndrome extraction to achieve this performance [42]. Such an approach requires access to 3 code blocks within the hardware (one block for the data, and two blocks for the maximally entangled resource state).
There is, however, a more compact option similar to Steane syndrome extraction [51], but derived from Knill’s approach to syndrome extraction. The key observation is illustrated in Fig. 5, resulting in a circuit that essentially serializes the syndrome extracting teleportation into two 1-bit teleportations [52]—the only requirement being that one must be able to prepare the encoded |$\ket{0}$ and $\ket{+}$ states, apply an encoded $\rm CNOT$, and measure in the encoded $X$ and $Z$ eigenbases (and all of these operations are available in the Carbon code).
Figure 5: Syndrome information can be obtained by performing teleportation at the logical level, as described by Knill [42]. Taking the realization of the original teleportation circuit for two logical qubits encoded in a Carbon block, which required 3 encoded blocks (left), it is possible to rearrange commuting circuit components to arrive at a circuit that uses a sequence of two 1-bit teleportations [52] to extract syndrome information requiring only 2 encoded blocks at any given time (right) (see Appendix A).
@@ IV.2 Physical baselines
Within a single block, the simplest fault-tolerant operations available for the Carbon code are the ${\rm H}\otimes{\rm H}$, and a $\rm CNOT$–the $\rm H$ is implemented via transversal $\rm H$ application followed by a qubit permutation, while the $\rm CNOT$ is implemented via a qubit permutation [53, 54]. Since qubit permutations can be applied in the H2 system without interaction between qubits, the permutations do not require additional ancillas to be fault-tolerant.
Combining gate teleportation and permutation-based gates allows for logical circuits of the form depicted in Fig. 6 in a single round of error correction. The $\rm CNOT$s and $\rm SWAP$s of logical qubits within a block and their composition can be implemented by qubit permutations.
Figure 6: Logical circuit family that can be implemented in a modified single round of error correction with the $[[12,2,4]]$ Carbon code, assuming $(\alpha,\beta,\delta,\gamma)\in\{0,\pm\frac{\pi}{2},\pm\pi\}$. The logical single-qubit rotations require modified state preparation during syndrome extraction, and the order of $X$ and $Z$ rotations can be transposed by modifying the order of the 1-bit teleportations, but neither of these modification impact the ability to extract syndrome information. Any combination of gates in the dashed boxes can be implemented by a permutation of the physical qubits.
@@
The comparison between a sequence of two $\rm CNOT$s and a round of error correction is motivated by the proposal in Ref. 23. Instead of benchmarking the entire circuit family implied by Fig. 6, we choose to compare against circuits with two $\rm CNOT$s since these are the physical gates with highest error rates in the physical system. Although in a fully scalable sub-threshold setting all logical gates can be made to improve upon their physical counterparts (by going to sufficiently high distance), the first bar that an experimental demonstration must meet is the demonstration of some non-trivial logical circuit that improves over its physical counterpart, which is our aim here.
@@ IV.3 Compiler optimizations
Time spent shuttling and cooling ions can lead to significant memory error, especially in error correction circuits that tend to leave some qubits idling for extended times [12]. Effective memory errors in H2 were previously reported in Ref. 22 for certain types of circuits with random connectivity, but the circuits used in this study are highly structured, leading to more coherent accumulation of noise, which is especially difficult to model accurately. To mitigate this problem, for the Carbon code experiments, we (1) modified the optimization of transport operations in the compiler and (2) incorporated dynamical decoupling pulses into the compiled circuits.
@@ IV.4 Experimental results
We observe a gain of more than an order of magnitude for a single round of error correction, as evidenced by a logical error rate of $0.017\%_{-0.001\%}^{+0.050\%}$, while the two physical baselines yield error rates of $1.0\%\pm 0.1\%$ (for a sequence of two 1-bit teleportations) and $0.43\%\pm 0.05\%$ (a sequence of two $\rm CNOT$s). For two rounds of error correction, the logical error rates increase to $0.3\%\pm 0.1\%$, while the two physical baselines’ error rates increase to $1.7\%_{-0.1\%}^{+0.2\%}$ and $0.76\%_{-0.08\%}^{+0.09\%}$ respectively, which represents a logical gain of more than 2 over the physical baselines. For three rounds of error correction the logical error rate increases to $0.5\%_{-0.2\%}^{+0.3\%}$, which still represents a gain of more than 2 over the physical baselines of $2.5\%\pm 0.2\%$ and $1.2\%\pm 0.1\%$. Details for each experiment and associated statistical uncertainties can be found in Table 3, and the results are illustrated in Fig. 7.
Although the gains over the physical error rates appear to behave non-linearly, the trend for the observed error rate per round of error correction is consistent with the expected linear behavior for a small number of repetitions, as illustrated in Fig. 8. The maximum-likelihood fits yield error rates per round of the 1-bit teleportation and the $\rm CNOT$ baseline of $0.7\%\pm 0.1\%$ and $0.42\%\pm 0.02\%$ respectively, while the fit for each round of error correction points to $0.3\%\pm 0.1\%$—indicating that the differences in the error rates per round are not statistically significant, largely due to the uncertainty in the firt for error correction.
The apparent non-linear trend in the gain is an artifact of the different $y$-intercepts for the linear trends of each of the sets of experiments. For the physical experiments, the $y$-intercept is set by the physical state-preparation and measurement (SPAM) error rate. This is also the case for the logical circuits, but details of the circuit implementation leads to a negative $y$-intercept despite positive SPAM error rates. This can be explained by considering optimizations in scheduling of the operations in the experimental device. In particular, since memory and transport errors are the dominating source of error in the experiments (see Section II), we estimate the optimized compiled circuit to have roughly $12r-$ 1 periods of significant waiting and transport for $r$ rounds of error correction (due to details of the initial state preparation and final state measurement), so the logical error rate for $r$ rounds is estimated to be roughly
@@
The pre- and post-rejection rates also serve to inform whether the error correction circuits are behaving as expected. In particular we expect the pre- and post-rejection rates to be linear in the number of repetitions (to leading order), and the experimental data is highly consistent with that prediction (see Appendix C).
@@
1 round unencoded baseline (2 teleports) $40,000$ 401 $1.0\%_{-0.1\%}^{+0.1\%}$
unencoded baseline (2 $\rm CNOT$s) $80,000$ 345 $0.43\%_{-0.05\%}^{+0.05\%}$
encoded, pre- and post-selection $10,000$ $7,094$ $7,008$ $1,418$ 1 $0.017\%_{-0.001\%}^{+0.050\%}$ 25–58
2 rounds unencoded baseline (4 teleports) $20,000$ 347 $1.7\%_{-0.1\%}^{+0.2\%}$
unencoded baseline (4 $\rm CNOT$s) $40,000$ 305 $0.76\%_{-0.08\%}^{+0.09\%}$
encoded, pre- and post-selection $14,548$ $7,528$ $7,294$ $2,892$ 22 $0.3\%_{-0.1\%}^{+0.1\%}$ 2.5–5
3 rounds unencoded baseline (6 teleports) $20,000$ 504 $2.5\%_{-0.2\%}^{+0.2\%}$
unencoded baseline (6 $\rm CNOT$s) $40,000$ 477 $1.2\%_{-0.1\%}^{+0.1\%}$
encoded, pre- and post-selection $10,000$ $4,029$ $3,588$ $1,896$ 19 $0.5\%_{-0.2\%}^{+0.3\%}$ 2.4–5
@@
Figure 7: Observed error rate for circuits with 1 to 3 rounds of error correction with the $[[12,2,4]]$ Carbon code (green circles) and physical baselines (blue squares for pairs of 1-bit teleportations, and orange x’s for pairs of $\rm CNOT$s). Results are offset along the x-axis for clarity. Linear fits are obtained by maximum-likelihood estimation (see Appendix B for details).
Figure 8: Incremental error rate obtained from maximum likelihood linear fits to the data in Fig. 7 (see Appendix B for details).
@@ V Summary and Outlook
We have demonstrated several fault-tolerant circuits outperforming their physical counterparts in a state-of-the-art QCCD trapped-ion system. Using different codes and protocols, we demonstrated error rates 4.7 to 800 times lower than the physical error rates in Bell correlation experiments with the $[[7,1,3]]$ and the $[[12,2,4]]$ codes. Moreover, we demonstrated up to ten rounds of error correction for the $[[12,2,4]]$ quantum code by combining error correction and error detection, with logical error rates well below a physical baseline circuit of two $\rm CNOT$s per round. With these results, we have demonstrated that current quantum processors are already able to reduce error rates in small circuits through quantum error correction. Future work will focus on extending these results to a richer set of fault-tolerant logical gadgets, and to ultimately enable universal fault-tolerant quantum computation while continuing to lower the achievable logical error rates. A significant milestone will be to demonstrate a universal family of quantum circuits with logical error rates significantly below the corresponding physical error rates.
@@ Acknowledgements.
We thank the broader quantum teams at Microsoft Azure Quantum and Quantinuum for many helpful discussions. We thank Jeongwan Haah for insightful discussions and helpful feedback, and Jonhas Colina for help with developing the compiler. A special thanks to Z. Alam and J. Strabley for making this collaboration possible.
Appendix A Carbon code from Knill’s C4/C6 scheme
A.1 Carbon definition and background
In 2005, Knill proposed a fault tolerance scheme based on concatenation of four-qubit ($C_{4}$) and six-qubit ($C_{6}$) codes [43]. This scheme boasts a circuit noise threshold of $3\%$, arguably the highest known, but also incurs large space and time overheads. Here, we describe an adaptation of Knill’s proposal at the first level of concatenation that substantially reduces the space and time overheads. At this first level, the concatenation of $C_{4}$ into $C_{6}$ yields a $[[12,2,4]]$ code, which we nickname Carbon ($C_{12}$). The stabilizers of the code are given in Table 4.
0 1 2 3 4 5 6 7 8 9 10 11
$X$ $X$ $X$ $X$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
$Z$ $Z$ $Z$ $Z$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $X$ $X$ $X$ $X$cdot $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $Z$ $Z$ $Z$ $Z$cdot $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $X$ $X$ $X$ $X$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $Z$ $Z$ $Z$ $Z$
$X$ $X$cdot $\cdot$ $\cdot$ $X$cdot $X$ $X$cdot $\cdot$ $X$
$X$cdot $\cdot$ $X$ $X$ $X$cdot $\cdot$ $\cdot$ $X$cdot $X$
$Z$cdot $Z$cdot $\cdot$ $\cdot$ $Z$ $Z$ $Z$cdot $\cdot$ $Z$
$Z$cdot $\cdot$ $Z$ $Z$cdot $Z$cdot $\cdot$ $\cdot$ $Z$ $Z$
$X$cdot $\cdot$ $X$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $X$ $X$ $=X_{0}$
$Z$cdot $\cdot$ $Z$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $Z$cdot $Z$ $=Z_{0}$
$\cdot$ $X$cdot $X$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $X$ $X$cdot $=X_{1}$
$Z$ $Z$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $Z$ $Z$cdot $=Z_{1}$
Table 4: Carbon stabilizers.
The $C_{4}$ code and its subsystem variants have been widely studied [57, 58, 59, 42, 43, 23] and used in proposals for magic state distillation [60, 61, 62] as well as fault-tolerant computation with linear qubit arrays [63, 64]. The $C_{6}$ code has been proposed for use in magic state distillation [65], but is much less widely studied.
Knill used teleportation as a means of error correction, this scheme achieves a remarkably high threshold of $3\%$ error per gate. A substantial drawback to Knill’s proposal is the qubit resource overhead. To execute an algorithm with circuit volume of $10^{8}$ with a physical CNOT error rate of $0.1\%$, Knill estimates a qubit overhead of $10^{5}$ per logical gate. There are a few reasons for the hefty overhead. First, the family of $C_{4}/C_{6}$ codes have even distance. There are some errors that can be detected but not corrected. In those situations, the multiple copies of circuits may be required in order to successfully execute a gate with high probability. Second, a key element of Knill’s proposal is to perform error correction by teleportation. However, the circuitry that Knill used was not particularly efficient—Knill’s $C_{4}$ preparation circuit requires eight qubits and eight CNOTs, whereas fault tolerant preparation circuits with five qubits and five CNOTs are now known [23]. Additionally, Knill did not consider optimizations across levels of concatenation.
Our adaptation benefits from treating the Carbon code monolithically rather than as a concatenation—making the circuits and the number of required ancillas much smaller. We leverage several modern fault tolerance innovations developed since Knill’s original proposal. The resulting circuits for preparing $|00\rangle$ require $19$ qubits and $37$ CNOTs, whereas Knill’s preparation requires $144$ qubits and $216$ CNOTs  [^6] The number of qubits depends on the amount of parallelism. Here we report the number of qubits required in order to achieve maximum parallelism. Our circuits for Bell state preparation require as few as 36 qubits. The reduction in circuit size directly improves acceptance rates, which reduces the expected number of copies required for success and further reduces qubit overhead.
A.2 Fault tolerance with Carbon
[…]
We limit our focus here to only those operations necessary for the error correction and Bell state demonstrations described in the main text, primarily |$\ket{00}$ and $\ket{++}$ preparation. In all cases, circuits for Carbon are designed in order to meet the following two conditions.
  1. 1.
    A logical error occurs only if there are at least three circuit faults.
  2. 2.
    Post-rejection occurs only if there are at least two circuit faults.
[…]
A circuit for preparing logical |$\ket{00}$ is illustrated in Fig. 9. Logical $\ket{++}$ can be prepared by appending transversal Hadamard and approprate qubit relabeling. Therefore, the |$\ket{00}$ circuit is the foundation for all of our teleportation circuits.
Figure 9: Carbon |$\ket{00}$ preparation circuit. The circuit consists of three parts. First three $C_{4}$ blocks are prepared, indicated by the gray, blue and green qubit lines. Then CNOTs are applied transversally between $C_{4}$ blocks in order to encode into Carbon. Then three weight-four stabilizers are measured to check for high weight errors. The $C_{4}$ preparations include flags, two of which are then combined with subsequent stabilizer measurements in order to reduce circuit size. Following the CNOT gates, permutation $(0~{}3~{}1~{}2)$ is applied to the middle four qubits (blue) and permutation $(0~{}2~{}3~{}1)$ is applied to the bottom four qubits (green).
In the absence of errors the measurements in Fig. 9 should all be $+1$. Any $-1$ outcome indicates an error and results in pre-rejection of the state. That is, the state is thrown away prior to interaction with any other encoded blocks. Conditioned on $+1$ outcomes, the output is “fault tolerant” in the following sense. For any single fault in the circuit, the residual error has CSS weight at most one, modulo stabilizers. CSS weight is defined as maximum Hamming weight of the $X$ and $Z$ bitstrings in the usual symplectic representation. For any two faults in the circuit, the residual error has CSS weight at most two or the residual error has a Carbon syndrome distinct from the syndrome of any CSS weight-one error (so that it can be post-rejected).
This circuit was designed by adapting and combining several existing circuits in the literature. First, Knill’s circuit for preparing |$\ket{00}$ for $C_{4}/C_{6}$ was stripped of all error detection circuitry, leaving only the $C_{4}$ preparations and transversal CNOTs between $C_{4}$ blocks. Next, flags were added to the $C_{4}$ preparations in order to detect weight-two errors [68]. (There is just a single weight-two error per preparation.) Note that other similar $C_{4}$ preparation circuits would suffice, such as [23]. Then additional stabilizer measurements were added to detect weight-two errors from the transversal CNOTs. The $X^{\otimes 4}$ and $Z^{\otimes 4}$ measurements on the middle four qubits use a circuit proposed by Reichardt [69]. The remaining $Z^{\otimes 4}$ measurement is a flag circuit modified to improve the depth [68].
The flag qubits used to detect errors in the $C_{4}$ preparations midway through the preparation are noteworthy. An alternative strategy is first prepare |$\ket{00}$ with a unitary circuit, and postpone flag qubits and measurements until after the unitary. This strategy is initially appealing because it is relatively easy to find unitary preparation circuits and, given a unitary preparation, it is easy to find sets of stabilizers that detect correlated errors. However, using flags early on in the circuit has two benefits. First, they can detect high-weight but manageable errors before they become even higher weight. A single fault in the middle $C_{4}$ preparation, for example, can cause a weight-two error $X_{7,8}$. The flag catches this error before it can subsequently spread to $X_{7,8,11,12}$ when it would be harder to manage. Second, fault tolerant detection circuitry is much simpler to construct. The $C_{4}$ flag gadgets don’t themselves cause any high weight errors. Contrast that with the alternative of measuring a weight-six stabilizer at the end of the circuit, which might require several additional flags.
A few variants of Fig. 9 may be useful. Two of the flags in the circuit can be combined with stabilizer measurements in order to reduce circuit size and measurement count. This may be beneficial if mid-circuit measurements are expensive. Another variant is to postpone the stabilizer measurements, as shown in Fig. 10. This can be useful when using |$\ket{00}$ as a component for a larger state. This approach is used, for example, to prepare Bell states, as described below.
A.2.2 Bell states
Since CNOT is transversal for Carbon, a pair of Bell states across two blocks can be prepared in the obvious way by first fault tolerantly preparing $\ket{++}$ and |$\ket{00}$ as described above, and then applying transversal CNOTs across the two blocks. An alternative approach is illustrated in Fig. 10. This approach also applies transversal CNOTs across $\ket{++}$ and |$\ket{00}$ blocks. But now stabilizer measurements are postponed until after the transversal CNOTs. Doing so allows for detecting additional errors on the transversal CNOTs, which could be an advantage depending on noise model and architecture. It can also be an advantage when preparing other entangled resource states for gate teleportation. The detection circuitry is shown in greater detail in Fig. 11.
Figure 10: Carbon Bell state preparation circuit. This circuit prepares two Bell states across two Carbon blocks. One Carbon block is prepared as $\ket{++}$ and the other as |$\ket{00}$. Note that the weight-four measurements illustrated in Fig. 9 are not necessary here. Then, corresponding $C_{4}$ blocks are paired up, followed by transversal CNOTs across each pair. Following the transversal CNOTs, $X^{\otimes 4}$ is measured on the top block of each pair and $Z^{\otimes 4}$ is measured on the bottom block of each pair. If any of the measurements have value $-1$, then the state is discarded. The $X^{\otimes 4}$, $Z^{\otimes 4}$ measurement circuit is given in Fig. 11.
Figure 11: Simultaneous measurement of $X^{\otimes 4}$ and $Z^{\otimes 4}$ on a pair of $C_{4}$ blocks. This circuit is similar to that of [63], but measures operators on two different code blocks. This circuit is fault tolerant in the sense that any single fault is either detected (by $-1$ outcome of a measurement) or results in an error of at most weight one (modulo $X^{\otimes 4}$ and $Z^{\otimes 4}$ on the data qubits. CNOTs on data qubits are shown sequentially for illustration, but may be re-ordered arbitrarily, or parallelized as desired.
A.2.3 Error correction
Given a Bell state, as prepared above, fault tolerant error correction can be implemented by teleportation, as described by Knill. Half of the Bell states, one of the Carbon blocks, interacts with an input Carbon “data” block. Then both the data and the Bell state half are measured. This has the effect of teleporting the data block into the remaining half of the Bell state and measuring a Carbon syndrome. The syndrome can then be used to detect or correct errors on the (teleported) data.
The syndrome can be decoded with a specially modified lookup decoder. A typical lookup decoder for an $n$-qubit code is constructed by enumerating over all $n$-qubit Pauli errors and computing each syndrome. But Carbon is a distance four code. Some distinct weight-two Pauli errors will generate the same syndrome and are therefore not correctable. To achieve an effective distance of five, so that logical errors occur only if there are at least three circuit faults, the lookup decoder can be modified so that uncorrectable syndromes are removed from the table. If such a syndrome occurs at runtime, then we “post-reject” by throwing those computations away.
In [43], Knill uses a phenomenological error model, in which errors occur independently on qubits, in order to determine which syndromes to remove from the lookup table. That model is not sufficient for our circuits. We allow pairs of faults to pass pre-selection if the syndrome of the residual error is distinct from that of all CSS-weight one Pauli errors. This yields efficient circuits, but can also allow for what appear to be correctable weight-two errors based on syndrome, but are actually errors of higher weight.
One way to account for this is by tailoring the lookup table according to the state preparation circuits for |$\ket{00}$ and $\ket{++}$. The table is constructed by enumerating over all single and double faults within the circuits. If, for a given syndrome, the residual error of the faults always matches the naive lookup table, then the syndrome is kept in the table. Otherwise, the syndrome is thrown out.
We, however, adopt a simpler stategy: start with the naive Carbon lookup decoder, but keep only the syndromes that are consistent with CSS weight-one Pauli errors. This option is more aggressive in that it throws away some error configurations that might be correctable, but it maintains a $O(p^{2})$ post-rejection rate. We find that it impacts the post-rejection rate only marginally, while substantially improving logical error rates.
A.3 Scaling to larger sizes
In Knill’s original proposal Carbon ($C_{4}/C_{6}$) was concatenated repeatedly with $C_{6}$ in order to increase the code distance. Knill suggested concatenating with other codes once the effective error rate is sufficiently low. The same approach is compatible with our scheme for Carbon. Our circuit optimizations do not immediately apply to further concatenations to $C_{6}$, as they do not respect the boundaries between $C_{4}$ and $C_{6}$, but Knill’s circuit constructions for $C_{6}$ can be applied without modification. Swapping out Knill’s $C_{6}$ encoding circuits for a flag-based stabilizer measurement approach is likely not beneficial because it would require selective CNOT gates, which are somewhat complex for these codes (require specialized state preparation and teleportation). Other optimizations may still be possible. For example, Knill uses Bell pairs followed by lower level error correction to encode $C_{6}$. When the lower level is Carbon, the error correction can be elided in favor of using Fig. 10. This saves a factor of three in qubit overhead.
Concatenation has the benefit of blunting the impact of post-selection. A syndrome that induces post-rejection in Carbon can instead be corrected as an erasure at higher levels of encoding. If error correction by teleportation is also used at higher levels, then Carbon gate gadgets will be entirely (or almost entirely) contained in state preparation circuits for those higher levels. Post-selection in Carbon is thereby converted into pre-selection for the larger code.
Appendix B Statistical analysis and error bars